OpenStudy (anonymous):

Find an equation of the tangent line to the curve y=6^x at the point (1,6) Tangent line : y = ? working on this a while please help!

zepdrix (zepdrix):

The equation of the line tangent to the curve, \(\large y=6^x\), at the point \(\large x=1\) will be of the form \(\large y=mx+b\) in slope-intercept form. Err actually, let's put it in point-slope form, that will a little bit less work. :) \[\large y-y_o=m(x-x_o)\] Where \(\large (x_o,y_o)\) is some point on the line. In this case our point is \(\large (1,6)\). \[\large y-6=m(x-1)\] The only thing we really have to do is ~ find m. \(\large m=f'(1)\). m is the slope of the tangent line at the point \(\large x=1\)

zepdrix (zepdrix):

So all we need to do is, take the derivative of our function, then plug \(\large x=1\) into it! Do you know how to take the derivative of an exponential function?

OpenStudy (anonymous):

got it thank you :)

zepdrix (zepdrix):

cool c: